A balloon ascends with a uniform acceleration of 981 /8 cm/s2, at the end of half a minute a body is released from it, find the time that elapses before the body reaches the ground. [g = 9.81]

Grade:11

1 Answers

vikas askiitian expert
13 years ago

accelration of baloon is 981/8 cm/sec^2=9.8/8 m/sec^2

after half minute velocity of baloon is v=u+at

                                                       =9.8*30/8 m/sec                 (u=0)

total height attained by baloon is H=ut+at^2/2

                                                  =450*9.8/8 m

 at this time the particle is dropped so its velocity will be same as of baloon in upward direction....

     velocity of paticle is u=9.8*30/8 m/sec

    now using s=ut+at^2/2 for particle

       s=-H , u=9.8*30/8 ,g=-9.8,after putting these values we get a quadratic in t

2t^2 -15t-225=0

 t=15,-7.5 sec

 t=15 is required ans